OpenStudy (anonymous):

the area of a rectangular wall of a barn is 36 square feet. Its length is 6 feet longer than twice its width. Find the length and width of the wall of the barn.

OpenStudy (anonymous):

area=lw l=6+2w substitute (6+2w)*w=36 6w+2w^2=36 2w^2+6w-36=0 2(w^2+3w-18)=0 2(w+6)(w-3)=0 w=3 or w=-6 width cannot be negative thus it must be 3 if the width is 3 then 36=3l 36/3=l 12=l

OpenStudy (anonymous):

good?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

np